40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11607     Entry time: Wed Sep 16 23:07:06 2015
Author: rana 
Type: Update 
Category: Electronics 
Subject: LSC Whitening board: LP filters added, pictures taken 

I added the 0.1 uF and 47 nF caps that I mentioned so that we can now bypass the AA filters for these channels. (mistakenly installed 47 instead of 0.47 nF on the first round and we got 350 Hz poles instead of 35 kHz)

Gautam and I checked out the AA sit and it seems that the XYCOM-220 cable which ought to allow switching of the AA filter is not connected on the XYCOM side, so the LSC AA filters are always ON. In order to bypass them, we'll need to just short the bypass control pins or just set the +5V on the board to GND, by lifting the EMI3 filter and shorting C6.

I have so far only made the changes on s/n 115 (used for AS55, REFL55, and REFL165), other 2 boards to follow soon.

Before making the AA change, we want to measure the HF spectrum the ADC for each of our main signals in the PRFPMI state. In lieu of that, we'll measure the spectrum at the I/Q mon ports of the demod boards via SR785 and then use matlab to propagate the signals to the ADC to make our estimate of how much anti-aliasing we need.

Changes relative to D990694-B:

  1. R215, R216, R217, R218, R219: 4.75k -> 9.53k.  This change was made long to make the DC gain of channels 4-8 be unity, the same as channels 1-3.
  2. 0.1 uF NPO cap in parallel with R127, R128, R129, R130, R131, R132, R133, R134.
  3. R127, R128, R129, R130, R131, R132, R133, R134 all 100k (was already like this) to keep LT1128 from floating up when input cables are disconnected.
  4. C158, C159, C160, C161, C162, C163, C164, C165, C166, C167, C168, C169, C170, C171, C172, C173, all were empty, now are 0.47 nF NPO.

I also looked at the noise in a few different configurations to see what we ought to do next.

BLACK: AS55I_IN1 with 0 dB whitening gain and whitening filter OFF, so its all just ADC noise

RED: same but with +45 dB whitening gain and WF ON, so above 10 Hz this is now the noise of the PD / demod chain

BLUE: RED w/ the anti-WF applied

PURPLE: in-loop POX11_I spectrum with x-arm locked

The conversion from counts to volts 0.0006, so the black trace is ~5 uV/rHz as expected. Its clear that we would be sort of OK for most of our channels if we just had 1 stage of whitening. I think we ought to convert the input stage into a 100:20 stage and also change the other whitenings into a 100:20 instead of 150:15. Then we'll have less gain at 15 Hz, but more at 100 Hz.

We really need to buy some surface mount capacitors, Steve - we ought to have at least 100 of all the ones in that little gray cabinet.

Attachment 1: 20150916_221210.jpg  2.137 MB  | Hide | Hide all
20150916_221210.jpg
Attachment 2: out.pdf  71 kB  Uploaded Thu Sep 17 00:48:33 2015  | Hide | Hide all
out.pdf
ELOG V3.1.3-